# DIPLODOCUS I: Framework for the evaluation of relativistic transport equations with continuous forcing and discrete particle interactions

### v1

@@ I Introduction
An alternate microscopic approach is to assume symmetries in the particle distribution in space and momentum space. For example, an emitting region of an astrophysical source can be simplified to a spherical or slab-like zone or zones with an internally isotropic distribution of particles. Early astrophysical use of these methods was in modelling cosmic ray transport [12, 13, 14]. Here, in addition to spherical symmetry in momentum space, specifically with reference to a co-moving frame, particle distributions were taken to be close to Maxwellian. The assumption of isotropy in particular led to extensive work [15, 16, 17, 18, 19, 20, 21, 22, 23] describing a wide range of interactions between particles and fields, with only small ventures into the world of anisotropy [24, 25]. These works form the foundation of modern single/multi-zone models for jetted astrophysical sources [26, 27, 28, 29, 30], which maintain the tradition of momentum-space isotropy to limit computational requirements. These requirements are further reduced by considering additional symmetries in the spatial distribution of particles. In general, such models operate by considering independent, homogenous zones of isotropic plasma, normally spherical or slab-like in shape, with prescribed internal properties such as bulk Lorentz factor, magnetic field strength and initial particle distributions. These zones may then placed along the axis of the jet to generate is structure, though typically zones do not interact with one another. Within these zones, the particle distributions are evolved according to their interactions and the prescribed parameters of the zone. Emission spectra are then generated by combining and Doppler boosting photon populations from each zone into an observer frame. These spectra can then be fit to observed data of astrophysical sources. These zoned model fits have provided great insight into the environment and dynamics of sources like AGN jets, allowing the study of questions such as how are particles accelerated to the initial distributions, and whether the jet material dominated by leptonic or hadronic particles, which are still of some debate [31, 32, 33].
@@ III.1 Phase Flow with Coordinates
As discussed in Section˜II.2, for a system to be obey Liouville’s theorem $\bm{\mathrm{d}}\bm{\omega}\overset{!}{=}0$, by consideration of Eq.˜24 this condition is equivalent to $\frac{\partial F^{a}}{\partial p^{a}}\overset{!}{=}0$. As an example where this condition is met, the Lorentz force on a particle of charge $q$ and mass $m$ is given by $F^{a}=\frac{q}{m}F^{ab}p_{b}$, where $F^{ab}$ is the electromagnetic field tensor, therefore $\frac{\partial F^{a}}{\partial p^{a}}=F^{a}_{~a}=0$ as $F^{ab}$ is antisymmetric. However, a counter example is the radiation reaction force [^7] The radiation reaction force is also known as the Abraham-Lorentz-Dirac force [66, 67, 68] [see also 69, 70, for more modern descriptions]. experienced by a charged particle being accelerated by an electromagnetic field where $F^{a}\propto\left(\eta^{ab}+\frac{p^{a}p^{b}}{m^{2}c^{2}}\right)\left(\frac{q}{m}p^{d}p^{c}\partial_{c}F_{bd}+\frac{q^{2}}{m^{2}}qF_{bc}F^{cd}p_{d}\right)$, for such a force it can be shown that $\frac{\partial F^{a}}{\partial p^{a}}\neq 0$, and is therefore a non-conservative force, evident by the fact that it results in the charged particle losing energy over time.
@@ IV Distribution-In-Plateaux
$\begin{split}H_{\alpha}(t)&=\Theta(t-t_{\alpha})-\Theta(t-t_{\alpha+1})\\ &=\begin{cases}1,&\text{for }t_{\alpha}<t<t_{\alpha+1},\\ h_{t}^{-}&\text{for }t=t_{\alpha}\\ h_{t}^{+}=1-h_{t}^{-}&\text{for }t=t_{\alpha+1}\\ 0,&\text{for all other }t,\end{cases}\end{split}$ (38)
with $\Theta$ being the Heaviside step function and $h_{t}^{-}$ is its value at zero for the coordinate tt [^9] By allowing the definition of $\Theta(0)$ to be a variable between 0 and 1, usually taken to be 0,1 or $1/2$, allows for slight modification of the numerical scheme, as described in Paper II, while maintaining conservative particle transport..
@@ IV.2 DIP Applied to Transport
With $\bm{n}$ and $e_{a}^{~\alpha}$ defined, the integral form of the transport equation Eq.˜10 can be evaluated using DIP over a phase space sub-domain $X_{\alpha\beta\gamma\delta}\times P_{ijk}=[t_{\alpha},t_{\alpha+1}]\times[x_{\beta},x_{\beta+1}]\times[y_{\gamma},y_{\gamma+1}]\times[z_{\delta},z_{\delta+1}]\times[p_{i},p_{i+1}]\times[u_{j},u_{j+1}]\times[\phi_{k},\phi_{k+1}]$ to give the rather long expressions Eqs.˜60, 61, 62, 63, 64, 66, 67, 68, 65, 69 and 71, whereby terms on the left hand side of Eq.˜10 are converted into fluxes at the boundaries of the sub-domain (see Fig.˜7) and the right hand side has been converted into a series of collision arrays. These equations form a numerical evolution scheme for all particle distribution functions, including external forces and collisions between particles. Such a scheme is conservative in terms of the total number of particles in the system (in the absence of emissive processes which are free to add or subtract particles from the system), as fluxes are always balanced between sub-domains, i.e. any particles leaving one sub-domain of phase space enter another.

### v2

@@ I Introduction
An alternate microscopic approach is to assume symmetries in the particle distribution in space and momentum space. For example, an emitting region of an astrophysical source can be simplified to a spherical or slab-like zone or zones with an internally isotropic distribution of particles. Early astrophysical use of these methods was in modelling cosmic ray transport [12, 13, 14]. Here, in addition to spherical symmetry in momentum space, specifically with reference to a co-moving frame, particle distributions were taken to be close to Maxwellian. The assumption of isotropy in particular led to extensive work [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] describing a wide range of interactions between particles and fields, with only small ventures into the world of anisotropy for specific [26, 27, 28, 29] and general [30] interactions. These works form the foundation of modern single/multi-zone models for jetted astrophysical sources [31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45], which maintain the tradition of momentum-space isotropy to limit computational requirements. These requirements are further reduced by considering additional symmetries in the spatial distribution of particles. In general, such models operate by considering independent, homogenous zones of isotropic plasma, normally spherical or slab-like in shape, with prescribed internal properties such as bulk Lorentz factor, magnetic field strength and initial particle distributions. These zones may then placed along the axis of the jet to generate is structure, though typically zones do not interact with one another. Within these zones, the particle distributions are evolved according to their interactions and the prescribed parameters of the zone. Emission spectra are then generated by combining and Doppler boosting photon populations from each zone into an observer frame. These spectra can then be fit to observed data of astrophysical sources. These zoned model fits have provided great insight into the environment and dynamics of sources like AGN jets, allowing the study of questions such as how are particles accelerated to the initial distributions, and whether the jet material dominated by leptonic or hadronic particles, which are still of some debate [46, 47, 48].
@@ III.1 Phase Flow with Coordinates
As discussed in Section˜II.2, for a system to be obey Liouville’s theorem $\bm{\mathrm{d}}\bm{\omega}\overset{!}{=}0$, by consideration of Eq.˜24 this condition is equivalent to $\frac{\partial F^{a}}{\partial p^{a}}\overset{!}{=}0$. As an example where this condition is met, the Lorentz force on a particle of charge $q$ and mass $m$ is given by $F^{a}=\frac{q}{m}F^{ab}p_{b}$, where $F^{ab}$ is the electromagnetic field tensor, therefore $\frac{\partial F^{a}}{\partial p^{a}}=F^{a}_{~a}=0$ as $F^{ab}$ is antisymmetric. However, a counter example is the radiation reaction force [^7] The radiation reaction force is also known as the Abraham-Lorentz-Dirac force [80, 81, 82] [see also 83, 84, for more modern descriptions]. experienced by a charged particle being accelerated by an electromagnetic field where $F^{a}\propto\left(\eta^{ab}+\frac{p^{a}p^{b}}{m^{2}c^{2}}\right)\left(\frac{q}{m}p^{d}p^{c}\partial_{c}F_{bd}+\frac{q^{2}}{m^{2}}qF_{bc}F^{cd}p_{d}\right)$, for such a force it can be shown that $\frac{\partial F^{a}}{\partial p^{a}}\neq 0$, and is therefore a non-conservative force, evident by the fact that it results in the charged particle losing energy over time.
@@ IV Distribution-In-Plateaux
$\begin{split}H_{\alpha}(t)&=\Theta(t-t_{\alpha})-\Theta(t-t_{\alpha+1})\\ &=\begin{cases}1,&\text{for }t_{\alpha}<t<t_{\alpha+1},\\ h_{\alpha}^{-}&\text{for }t=t_{\alpha}\\ h_{\alpha}^{+}=1-h_{\alpha}^{-}&\text{for }t=t_{\alpha+1}\\ 0,&\text{for all other }t,\end{cases}\end{split}$ (38)
with $\Theta$ being the Heaviside step function and $h_{\alpha}^{\pm}$ being the boxcar functions values at the bounds of the domain $t_{\alpha}\leq t\leq t_{\alpha+1}$ [^9] By allowing the definition of $\Theta(0)$ to be a variable between 0 and 1, usually taken to be 0,1 or $1/2$, modification of the numerical scheme may be made, as described in Paper II, while maintaining conservative particle transport..
@@ IV.2 DIP Applied to Transport
With $\bm{n}$ and $e_{a}^{~\alpha}$ defined, the integral form of the transport equation Eq.˜10 can be evaluated using DIP over a phase space sub-domain $X_{\alpha\beta\gamma\delta}\times P_{ijk}=[t_{\alpha},t_{\alpha+1}]\times[x_{\beta},x_{\beta+1}]\times[y_{\gamma},y_{\gamma+1}]\times[z_{\delta},z_{\delta+1}]\times[p_{i},p_{i+1}]\times[u_{j},u_{j+1}]\times[\phi_{k},\phi_{k+1}]$ to give the rather long expressions Eqs.˜60, 61, 62, 63, 64, 66, 67, 68, 65, 69 and 71, found in Appendix˜C, whereby terms on the left hand side of Eq.˜10 are converted into fluxes at the boundaries of the sub-domain (see Fig.˜7) and the right hand side has been converted into a series of collision arrays. These equations form a numerical evolution scheme for all particle distribution functions, including external forces and collisions between particles. Such a scheme is conservative in terms of the total number of particles in the system (in the absence of emissive processes which are free to add or subtract particles from the system), as fluxes are always balanced between sub-domains, i.e. any particles leaving one sub-domain of phase space enter another.